Two balls are drawn in succession without replacement from an urn containing 5 white balls and 6 black balls. Let Z be the random variable representing the number of black balls. Construct the probability distribution of the random variable Z.

Possible Outcomes-


Value of the Random Variable Z(Number of Blue Balls)-


1
Expert's answer
2021-05-24T16:13:55-0400
"P(WW)=\dfrac{5}{11}(\dfrac{4}{10})=\dfrac{2}{11}"

"P(WB)=\dfrac{5}{11}(\dfrac{6}{10})=\dfrac{3}{11}"

"P(BW)=\dfrac{6}{11}(\dfrac{5}{10})=\dfrac{3}{11}"

"P(BB)=\dfrac{6}{11}(\dfrac{5}{10})=\dfrac{3}{11}"

Let "Z=" Number of Blue Balls. Then


"\begin{matrix}
 z & 0 & 1 & 2 \\
\\
 p(z) & \dfrac{2}{11} & \dfrac{6}{11} & \dfrac{3}{11}
\end{matrix}"

Check

"\dfrac{2}{11}+\dfrac{6}{11}+\dfrac{3}{11}=1"